# Berezinskii-Kosterlitz-Thouless quantum transition in 2 dimensions

### v1

@@ Abstract
A continuous quantum phase transition in D dimensions is governed by the action of a field theory in (D+1) Euclidean dimensions (for a review see negele ). It is generally believed that there cannot be a quantum ($T=0$) BKT transition driven by a coupling constant in a 2D model because the dimension of the Euclidean action would be 3, and would thus be too high (see, e.g. feigelman ). In this note we show that this is wrong. There are indeed 2D models that undergo a quantum BKT transition at $T=0$ and these involve gauge fields with a diverging dielectric constant.
@@
Intuitively it is clear that a diverging dielectric constant $\varepsilon$ implies a dimension less, like infinite temperature. Recalling that the velocity of light in a medium is given by $v=1/\sqrt{\varepsilon\mu}$ (we use natural units $c=1$, $\hbar=1$ and $\varepsilon_{0}=1$), with $\mu$ the magnetic permeability (which we will henceforth consider to be $\mu=1$ for simplicity), we see that the divergence of $\varepsilon$ implies a vanishing light velocity $v=0$. Therefore, nothing propagates in the medium, only static electric configurations matter, as follows from the 2D Maxwell equation
@@
where $f^{2}$ is the usual massive coupling of (2+1)-dimensional electromagnetism which, in the case of applications to thin films, is given by $f^{2}=d/e^{2}$ with $d$ the film thickness and $e$ the electron charge. The corresponding Lagrangian is
@@
from where it is evident that magnetic fields become irrelevant in the limit $v\to 0$, $f^{2}\to\infty$ with $v^{2}f^{2}=\phi^{2}={\rm constant}$, which corresponds to the strong coupling regime of the model. In this limit, the Lagrangian becomes positive definite and involves only static electric field configurations.
@@
where now ${\bf E}=-\nabla vA_{0}-\partial_{\tau}{\bf A}$. Finally, since electric fields are static, we can evaluate the integral over $\tau$ to obtain
@@
Thin films are typically modelled as systems in which the matter degrees of freedom are constrained to live on a 2D plane, while electromagnetic films can space out in 3D. But do they really do so? Let us consider the experimentally relevant configuration of a horizontal thin film of large dielectric constant $\varepsilon$. In this geometric arrangement, there is necessarily a half-space above the film and one below the film and these can be considered as filled by media, which can of course also be the vacuum. So, above the film there is a medium of dielectric constant $\varepsilon_{1}$ and below the film there is a medium of dielectric constant $\varepsilon_{2}$. If the film is suspended in vacuum, we have $\varepsilon_{1}=\varepsilon_{2}=1$. If the film is deposited on a substrate, only $\varepsilon_{1}=1$. We are interested in the case in which $\varepsilon_{1}\ll\varepsilon$ and $\varepsilon_{2}\ll\varepsilon$. This problem has been considered and solved in chaplik ; keldysh (for a review see baturina ) and leads to the electric Keldish potential, which is a purely 2D logarithmic Coulomb potential up to an electric screening length $\lambda_{s}=d\varepsilon/(\varepsilon_{1}+\varepsilon_{2})$, with $d$ the film thickness.
This result can also be obtained with a much simpler calculation in the symmetric situation where the media above and below the film are the same. In this case, the behaviour of the electric field at the upper and lower boundaries of the film must be the same by symmetry, which implies periodic boundary conditions. With periodic boundary conditions, the Coulomb potential $G$ among two charges $e$ in a film of thickness $d$ along the z-axis is determined by the equation
@@
In the regime $d\ll|{\bf x}|\ll\varepsilon d$ we can keep only the first dipole $n=\pm 1$ and neglect $z$ in the ensuing cosine function, which gives
@@
up to a screening length $\lambda_{s}$. Up to a numerical constant O(1) this is the Keldish result. The Keldish potential implies that, for very large dielectric constant $\varepsilon$, the electric screening length becomes typically larger than the lateral dimension $L$ of the sample, $L<\lambda_{s}$. In these films, electromagnetism is entirely 2D: the lines of electric field do not exit the film in the orthogonal direction.
This has a momentous consequence if the material in the film is a superconductor. Indeed, the time-dependent Ginzburg-Landau theory becomes a super-renormalizable, purely 2D theory and, when the strong electric fluctuations are taken into account, is therefore plagued by infrared divergences. These can be cured by an expansion in terms of ${\rm ln}(e^{2}L/d)$ jackiw , which means the theory becomes non-perturbative in systems larger than their thickness. When decreasing the thickness, a film of a superconductor with a large normal-state dielectric constant breaks thus up in many condensate islands of the typical lateral size $d$, each characterized by its own phase, forming an emergent Josephson junction array (JJA) super2D ; typeIII ; typeIIIvort . This has been verified for such materials as TiN sacepe1 , NbTiN nbtin and InO granular ; kapitulnik . It becomes now clear that, in these granular materials, it is the compact version of electrodynamics which is the relevant one. Each condensate island is characterized by its own independent phase: since this is an angular variable, the gauge group is the compact group $U(1)$. The compactness of the gauge group implies the presence of topological excitations absent in the non-compact version of electrodynamics. We must thus amend the above computation of the effects of a large dielectric constant to take these topological excitations into account.
@@
where $a_{\mu}$ is the emergent gauge field, $\bf e$} and $b$ the corresponding emergent electric and magnetic field, respectively, $\varphi^{\mu}$ the integer vortex three-current and $k^{2}$ the relevant mass coupling constant. In the limit of high effective dielectric constant there are no currents and again, only static emergent electric configurations matter. We can repeat verbatim the steps above to obtain the action
@@
For $\eta>1$, we have thus two quantum BKT transitions at $g=\eta$ and at $g=1/\eta$. For high values of $g>\eta$ we have a superconductor, for very low values of $g<1/\eta$ we have a superinsulator. At both transitions the effective dielectric constant diverges. In between there is a Bose metal phase, a topological state in which both charges and vortices are out of condensate but frozen by statistical interactions dst ; bm (for a review see boserev ). For $\eta<1$, instead, there is a domain $\eta<g<1/\eta$ in which superinsulation can coexist with superconductivity, a superinsulator being the most stable state for $g<1$ while a superconductor is favoured for $g>1$. This is essentially a “first-order quantum transition” in which the changeover from superconductivity to superinsulation is discontinuous dst ; bm . The two lines of phase transitions between superconductor and Bose metal and between Bose metal and superinsulator and the interactions between charges in these three phases have been derived analytically in dst ; bm ; dtv . Here the focus is on their BKT universality class and the generic mechanism by which quantum BKT transitions in 2D may take place.
@@
Of particular interest is the extreme case $z\to\infty$, which is normally associated with a quantum Griffiths transition (for a review see vojta ) corresponding to “infinite” disorder. Our result shows that this is not necessarily the case. As we have shown, the extreme non-relativistic case corresponding to $z\to\infty$ is realized as well for quantum BKT transitions due to a diverging dielectric constant, in which case it has nothing to do with disorder and can occur also in perfectly ordered JJA. This is what has induced some authors xing ; zhang ; jing ; yadav to associate the superconductor to Bose metal transition with a quantum Griffiths transition due to extreme disorder. This is already an improbable scenario because it would entail a discontinuity of the quantum transition at $T=0$ with respect to $T>0$, where it is known to be a BKT transition. Given that, in these thin superconducting films, the dielectric constant is known to diverge at this transition, the alternative scenario of a quantum BKT transition independent of disorder bm should be highly favoured.

### v2

@@ I Introduction
The original BKT transition is driven by the condensation of vortices as a function of temperature. Here we shall consider a different BKT transition, governed by a coupling constant, rather than temperature, but still in 2 space dimensions (2D). A continuous quantum phase transition in D dimensions is governed by the action of a field theory in (D+1) Euclidean dimensions (for a review see negele ). It is often stated that there cannot be a quantum ($T=0$) BKT transition driven by a coupling constant in a 2D model because the dimension of the Euclidean action would be 3, and would thus be too high (see, e.g. feigelman ). In this note we show that this is wrong. There are indeed 2D models that undergo a quantum BKT transition at $T=0$ and these involve gauge fields with a diverging dielectric constant. The mechanism is the same as in the original BKT transition but it is driven by a different class of topological defects.
@@
Intuitively, it is clear that a diverging dielectric constant $\varepsilon$ implies a dimension less, like infinite temperature. Recalling that the velocity of light in a medium is given by $v=1/\sqrt{\varepsilon\mu}$ (we use natural units $c=1$, $\hbar=1$ and $\varepsilon_{0}=1$), with $\mu$ the magnetic permeability (which we will henceforth consider to be $\mu=1$ for simplicity), we see that the divergence of $\varepsilon$ implies a vanishing light velocity $v=0$. Therefore, nothing propagates in the medium, only static electric configurations matter, as follows from the 2D Maxwell equation
@@
where $f^{2}$ is the usual massive coupling of (2+1)-dimensional electromagnetism which, in the case of applications to thin films, is given by $f^{2}=e^{2}/d$ with $d$ the film thickness and $e$ the electron charge. The corresponding Lagrangian is
@@
from where it is evident that magnetic fields become irrelevant in the limit $v\to 0$, $f^{2}\to\infty$ with $v^{2}f^{2}=\phi^{2}={\rm constant}$, which corresponds to the strong coupling regime of the model. In this limit the Lagrangian becomes
$L={1\over 2\phi^{2}}\int d^{2}x\ {\bf E}^{2}\ ,$ (4)
which is positive definite and involves only static electric field configurations.
@@
where now ${\bf E}=-\nabla(A_{0}/v)-\partial_{\tau}{\bf A}$. Finally, since electric fields are static, we can evaluate the integral over $\tau$ to obtain
@@
II The Keldysh potential
Thin films are typically modelled as systems in which the matter degrees of freedom are constrained to live on a 2D plane, while electromagnetic fields can space out in 3D. But do they really do so? Let us consider the experimentally relevant configuration of a horizontal thin film of large dielectric constant $\varepsilon$. In this geometric arrangement, there is necessarily a half-space above the film and one below the film and these can be considered as filled by media, which can of course also be the vacuum. So, above the film there is a medium of dielectric constant $\varepsilon_{1}$ and below the film there is a medium of dielectric constant $\varepsilon_{2}$. If the film is suspended in vacuum, we have $\varepsilon_{1}=\varepsilon_{2}=1$. If the film is deposited on a substrate, only $\varepsilon_{1}=1$. We are interested in the case in which $\varepsilon_{1}\ll\varepsilon$ and $\varepsilon_{2}\ll\varepsilon$. This problem has been considered and solved in chaplik ; keldysh (for a review see baturina ) and leads to the electric Keldish potential, which is a purely 2D logarithmic Coulomb potential up to an electric screening length $\lambda_{s}=d\varepsilon/(\varepsilon_{1}+\varepsilon_{2})$, with $d$ the film thickness.
This result can also be obtained with a much simpler calculation in the symmetric situation where the media above and below the film are the same. In this case, the behaviour of the electric field at the upper and lower boundaries of the film must be the same by symmetry: both normal components vanish due to the much higher dielectric constant in the film than outside and therefore the potential must also be the same since a normal component inside the film has nowhere to end. This implies periodic boundary conditions. With periodic boundary conditions, the Coulomb potential $G$ among two charges $e$ in a film of thickness $d$ along the z-axis is determined by the equation
@@ II The Keldysh potential
In the regime $|{\bf x}|\gg\varepsilon d$ the prefactor of $t$ in the phase is very large even for $n=\pm 1$. Therefore the phase factor oscillates wildly and the whole expression is suppressed by the integral over $t$. In the opposite regime $|{\bf x}|\ll\varepsilon d$ the phase again oscillates strongly for high values of $n$. In this regime, the dominant contribution that survives comes from the first dipole $n=\pm 1$. In this regime we can thus keep only the first dipole $n=\pm 1$ and neglect $z$ in the ensuing cosine function, which gives
@@
up to a screening length $\lambda_{s}$. Up to a numerical constant O(1) this is the Keldish result. The Keldish potential implies that, for very large dielectric constant $\varepsilon$, the electric screening length in very thin films can become typically larger than the lateral dimension $L$ of the sample, $L<\lambda_{s}$. In films where this happens, electromagnetism is entirely 2D: the lines of electric field do not exit the film in the orthogonal direction.
This has a momentous consequence if the material in the film is a superconductor. Indeed, the time-dependent Ginzburg-Landau theory becomes a super-renormalizable, purely 2D theory and, when the strong electric fluctuations are taken into account, is therefore plagued by infrared divergences. These can be cured by an expansion in terms of ${\rm ln}(e^{2}L/d)$ jackiw , which means the theory becomes non-perturbative in systems larger than their thickness. When decreasing the thickness, a film of a superconductor with a large normal-state dielectric constant breaks thus up in many condensate islands of the typical lateral size $d$, each characterized by its own phase, forming an emergent Josephson junction array (JJA) super2D ; typeIII ; typeIIIvort . This has been verified for such materials as TiN sacepe1 , NbTiN nbtin and InO granular ; kapitulnik . It becomes now clear that, in these granular materials, it is the compact version of electrodynamics which is the relevant one. Each condensate island is characterized by its own independent phase: since this is an angular variable, the gauge group is the compact group $U(1)$. The compactness of the gauge group implies the presence of topological excitations, absent in the non-compact version of electrodynamics. We must thus amend the above computation of the effects of a large dielectric constant to take these topological excitations into account.
III Compact QED with a diverging dielectric constant
@@
We would like to conclude this section with a caveat and a note. First of all, the existence of the 2D quantum BKT transition relies on the particular limit described above and it is not guaranteed that all systems will end up in this regime. Our point is to prove that such a quantum BKT transition in 2D can exist, which is totally new and unexpected. Secondly, its realization depends on an underlying structure and the effective electromagnetic action being the compact version. This is not exotic. As we have seen, the condensate of thin superconducting films near the SIT typically decomposes into grains that provide the necessary structure and the superinsulating state of such films does indeed realize the compact version of QED.
IV Vortex transition and the SIT
@@
where $a_{\mu}$ is the emergent gauge field, $\bf e$} and $b$ the corresponding emergent electric and magnetic fields, respectively, $\varphi^{\mu}$ the integer vortex three-current and $k^{2}$ the relevant mass coupling constant. In the limit of high effective dielectric constant there are no currents and again, only static emergent electric configurations matter. We can repeat verbatim the steps above to obtain the action
@@
For $\eta>1$, we have thus two quantum BKT transitions at $g=\eta$ and at $g=1/\eta$. For high values of $g>\eta$ we have a superconductor, for very low values of $g<1/\eta$ we have a superinsulator. At both transitions the effective dielectric constant diverges. In between there is a Bose metal phase, a topological state in which both charges and vortices are out of condensate but are frozen by statistical interactions dst ; bm (for a review see boserev ). For $\eta<1$, instead, there is a domain $\eta<g<1/\eta$ in which superinsulation can coexist with superconductivity, a superinsulator being the most stable state for $g<1$ while a superconductor is favoured for $g>1$. This is essentially a “first-order quantum transition” in which the changeover from superconductivity to superinsulation is discontinuous dst ; bm . The two lines of phase transitions between superconductor and Bose metal and between Bose metal and superinsulator and the interactions between charges in these three phases have been derived analytically in dst ; bm ; dtv . Here the focus is on their BKT universality class and the generic mechanism by which quantum BKT transitions in 2D may take place.
V Diverging dynamical exponent $z$ with no disorder
@@
Of particular interest is the extreme case $z\to\infty$, which is normally associated with a quantum Griffiths transition (for a review see vojta ) corresponding to “infinite” disorder. Our result shows that this is not necessarily the case. As we have shown, the extreme non-relativistic case corresponding to $z\to\infty$ is realized as well for quantum BKT transitions due to a diverging dielectric constant, in which case it has nothing to do with disorder and can occur also in perfectly ordered JJA. This is what has induced some authors xing ; zhang ; jing ; yadav to associate the superconductor to Bose metal transition with a quantum Griffiths transition due to extreme disorder. This is already an improbable scenario because it would entail a discontinuity of the quantum transition at $T=0$ with respect to $T>0$, where it is known to be a BKT transition. Given that, in these thin superconducting films, the dielectric constant is known to diverge at this transition, the alternative scenario of a quantum BKT transition independent of disorder bm should be highly favoured. Of course, in this case also the critical exponent $\nu$ diverges. In other words, a diverging $z$ exponent can be indicative of both a quantum Griffiths transition or a 2D quantum BKT transition. However, the Griffiths scenario must be discarded in favour of the 2D quantum BKT transition when the $z$ divergence occurs in a perfectly ordered system.